OpenStudy (anonymous):

Find the type and number of solutions for the equation. x^2 - 4x = -4 A)one distinct real solution B)two distinct real solutions C) two distinct nonreal complex solutions

zepdrix (zepdrix):

Write it in that nice quadratic form first of all! c: \[x^2 -4x + 4 = 0\] then lets see.. hmmmm

zepdrix (zepdrix):

looks like it is a nice perfect square, doesn't it? c: do ya see it?? :D

OpenStudy (anonymous):

yeh

OpenStudy (anonymous):

but what does that mean?

zepdrix (zepdrix):

the equation simplifies really nicely to\[(x-2)^2 = 0\] so hmm solving for x.. we can square both sides of this equation.

OpenStudy (anonymous):

so would it be A?

zepdrix (zepdrix):

ya that sounds right c: x=2 is the only solution

OpenStudy (anonymous):

(x-2)(x-2)=0 x=2 :)